PSNFSS: test most of the ‘35’ PostScript fonts

Sebastian Rahtz  Walter Schmidt

2000–01–12

This is typeset in in Adobe AvantGarde (pag):

A regular fée. An italic fée. A SmallCaps fée. A slanted fée. A bold fée. A bold italic fée. A bold SmallCaps fée. A bold slanted fée.

This is typeset in in Adobe Bookman (pbk):

A regular fée. An italic fée. A SmallCaps fée. A slanted fée. A bold fée. A bold italic fée. A bold SmallCaps fée. A bold slanted fée.

This is typeset in in Adobe Courier (pcr):
A regular fée. An italic fée.
A SmallCaps fée.
A slanted fée.
A bold fée. A bold italic fée.
A bold SmallCaps fée.
A bold slanted fée.

This is typeset in in Adobe Helvetica (phv):

A regular fée. An italic fée. A SmallCaps fée. A slanted fée.

A bold fée. A bold italic fée. A bold SmallCaps fée. A bold slanted fée.

This is typeset in in Adobe NewCenturySchoolbook (pnc):

A regular fée. An italic fée. A SmallCaps fée. A slanted fée. A bold fée. A bold italic fée. A bold SmallCaps fée. A bold slanted fée.

This is typeset in in Adobe Palatino (ppl):

A regular fée. An italic fée. A SmallCaps fée. A slanted fée. A bold fée. A bold italic fée. A bold SmallCaps fée. A bold slanted fée.

This is typeset in in Adobe Times Roman (ptm):

A regular fée. An italic fée. A SmallCaps fée. A slanted fée. A bold fée. A bold italic fée. A bold SmallCaps fée. A bold slanted fée.

This is typeset in Adobe ZapfChancery (pzc):

An italic fée.

In Helvetica we can try the narrow font series, too.^{1}

A narrow fée. A narrow slanted fée. A bold narrow fée. A bold narrow slanted fée.

1Thanks to Frank Poppe for putting this in.